# Nash CoT: Multi-Path Inference with Preference Equilibrium

### v1

@@ 7 Conclusion
In this study, we proved the existence of Nash equilibrium in preference model, subsequently, we proposed a new CoT approach Nash CoT, and validated its performance on various inference benchmarks. Experimental results show that Nash CoT can perform equally or even better than self-consistency while only require half inference costs.
@@
Despite Nash CoT showcase competitive performance with only half of inference paths, it requires pre-defined template, thus it’s in-convenient to utilize Nash CoT in new emerging scenario, in the future we will develop a automatic approach to balance task feedback and template design.
@@
Despite LLM has showcased superiority performance on vast benchmarks, but pre-train or fine-tune a LLM requires numerous computing resources. Therefore, it’s crucial to study how to inference a LLM to reach the ceiling of its capacity. CoT is a ideal approach which has been proved that can obviously evaluate the performance of LLMs’ inference. Among that,self-consistency is one of the best CoT approach.
Our method effectively reduce the inference times of multi-path inference, thereby reducing the deploy budgets of self-consistency. We believe our approach can further elevate the effectiveness of multi-path inference, thereby further improving the effectiveness of LLM.
acl
@@ Appendix A Dataset
Our majority dataset are composed of three different kinds of inference tasks. 1) arithmetic reasoning: SingleEq (Koncel-Kedziorski et al., 2015), AddSub (Hosseini et al., 2014), MultiArith (Roy and Roth, 2016), GSM8K (Cobbe et al., 2021), AQUA (Ling et al., 2017), and SVAMP (Patel et al., 2021). 2) symbolic reasoning: Last Letters, Coin Flip (Wei et al., 2023), and Object Tracking, Bigbench Date. 3) commonsense question answering: CommonsenseQA Talmor et al. (2019) and StrategyQA Geva et al. (2021). For more details about the dataset please refer to Wang et al. (2023).
Appendix B Uesage of LLM.
@@
Appendix C Computing Resources
@@
Appendix D Source Code.
@@
Appendix E Proof of theorem 3.1.
@@
where $\bar{\pi}$ and $\bar{\pi}^{\prime}$ are any two given policy set. Subsequently, we can further darrive at the following relationships:
$\begin{split}(v(\bar{\pi})-v(\bar{\pi}^{\prime}))^{\rm T}(\bar{\pi}-\bar{\pi}^{\prime})&=\begin{pmatrix}R(\pi_{1};\pi_{2})-R(\pi_{1}^{\prime};\pi_{2}^{\prime})\\ R(\pi_{2};\pi_{1})-R(\pi_{2}^{\prime};\pi_{1}^{\prime})\\ \end{pmatrix}\cdot\begin{pmatrix}\pi_{1}-\pi_{1}^{\prime},\pi_{2}-\pi_{2}^{\prime}\\ \end{pmatrix}\\ &=\bigg{(}R(\pi_{1};\pi_{2})-R(\pi_{1}^{\prime};\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime})+\bigg{(}R(\pi_{2};\pi_{1})-R(\pi_{2}^{\prime};\pi_{1}^{\prime})\bigg{)}\cdot(\pi_{2}-\pi_{2}^{\prime})\\ &=\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime})+\bigg{\{}2-\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})+\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\bigg{\}}\\ &\cdot(\pi_{2}-\pi_{2}^{\prime})\\ &=\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime}-\pi_{2}+\pi_{2}^{\prime})+2\cdot(\pi_{2}-\pi_{2}^{\prime})\\ &=\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})-2\bigg{)}\cdot(\pi_{2}^{\prime}-\pi_{2})+\\ &\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime}).\end{split}$ (6)

### v2

@@ 7 Conclusion
We prove the existence of a Nash equilibrium in preferences between role-immersed and normal-status LLM generations, and develop a gaming system within each path of the multi-path CoT to introduce Nash CoT. Experimental results show that Nash CoT can perform equally or even better than self-consistency while requiring fewer paths.
@@
Since Nash CoT relies on predefined templates, its performance will decline in new scenarios due to limited adaptability in template selection. Future efforts will concentrate on automating the balancing of task feedback and template selection.
Acknowledgement
This work was supported by the National Science and Technology Innovation 2030 - Major Project (Grant No.2022ZD0208800), and NSFC General Program (Grant No.62176215). Meanwhile, we thank the reviewers for their valuable suggestions.
@@ Acknowledgement
Pre-training or fine-tuning an LLM requires much more computing resources. Multi-path CoT is an ideal approach that has been proven to enhance the performance of LLMs’ inference. Meanwhile, Nash CoT effectively reduces the inference paths needed and times of multi-path inference. We believe our approach can further elevate the effectiveness of multi-path inference, thereby further improving the effectiveness of LLM.
@@ Appendix A Dataset
Our majority dataset are composed of three different kinds of inference tasks.
  • Arabic Reasoning: SingleEq (Koncel-Kedziorski et al., 2015), AddSub (Hosseini et al., 2014), MultiArith (Roy and Roth, 2016), GSM8K (Cobbe et al., 2021), AQUA (Ling et al., 2017), SVAMP (Patel et al., 2021), and MATH dataset (Hendrycks et al., 2021). In particular, MATH dataset are process by our own, : When constructing the dataset, we identify the answers to each question in the MATH dataset through ground truth (GT) reasoning, retaining a certain number of questions for each category. When testing the LLM, we input the questions, collect the responses from the deployed LLM, and determine whether the LLM’s reasoning is correct by comparing whether the inferred results match the GT, examples are shown in Table 11. Meanwhile, the dataset or its url will be released at our codebase page.
    Dataset Capacity
    Algebra 778
    Counting and Probability 291
    Geometry 235
    Intermediate Algebra 443
    Number Theory 474
    Prealgebra 618
    Precalculus 133
    Table 8: Data type and data capacity
  • Symbolic Reasoning: Last Letters, Coin Flip (Wei et al., 2023), and Object Tracking, Bigbench Date.
  • Commonsense Question Answering: CommonsenseQA Talmor et al. (2019) and StrategyQA Geva et al. (2021). For more details about the dataset please refer to Wang et al. (2023).
Appendix B Experimental Settings
In this section, we compare Nash CoT with self-consistency, focusing specifically on self-consistency as a representative method. As illustrated in Figure 1, Nash CoT operates with two distinct loops, where the final answer is derived from voting on the response that achieves Preference Equilibrium most frequently. In contrast, self-consistency determines the most frequent answer across all predictions.
Figure 5: Comparison of Nash CoT and self-consistency. (right) Nash Chain of thought (Nash CoT). (left) self-consistency.
Appendix C Usage of LLM.
@@ Appendix B Experimental Settings
Appendix D Computing Resources
@@
Appendix E Source Code.
@@
Appendix F Proof of theorem 3.1.
@@
where $\bar{\pi}$ and $\bar{\pi}^{\prime}$ are any two given policy set. Subsequently, we can further arrive at the following relationships:
$\begin{split}(v(\bar{\pi})-v(\bar{\pi}^{\prime}))^{\rm T}(\bar{\pi}-\bar{\pi}^{\prime})&=\begin{pmatrix}R(\pi_{1};\pi_{2})-R(\pi_{1}^{\prime};\pi_{2}^{\prime})\\ R(\pi_{2};\pi_{1})-R(\pi_{2}^{\prime};\pi_{1}^{\prime})\\ \end{pmatrix}\cdot\begin{pmatrix}\pi_{1}-\pi_{1}^{\prime},\pi_{2}-\pi_{2}^{\prime}\\ \end{pmatrix}\\ &=\bigg{(}R(\pi_{1};\pi_{2})-R(\pi_{1}^{\prime};\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime})+\bigg{(}R(\pi_{2};\pi_{1})-R(\pi_{2}^{\prime};\pi_{1}^{\prime})\bigg{)}\\ &\cdot(\pi_{2}-\pi_{2}^{\prime})\\ &=\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime})+\bigg{\{}2-\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})+\\ &\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\bigg{\}}\cdot(\pi_{2}-\pi_{2}^{\prime})\\ &=\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime}-\pi_{2}+\pi_{2}^{\prime})+2\cdot(\pi_{2}-\pi_{2}^{\prime})\\ &=\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})-2\bigg{)}\cdot(\pi_{2}^{\prime}-\pi_{2})+\\ &\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime}).\end{split}$ (6)
@@
Appendix G Further clarify our settings
Why in Table 12 and 3 do we set the number of pathways for Nash CoT to be half of self-consistency?
If Nash CoT can surpass self-consistency with half the number of reasoning pathways, then it is easier to demonstrate the validity and efficiency of Nash CoT. Because that according to the experimental conclusion in Wang et al. that ‘‘The experimental effect of self-consistency increases as the number of pathways increases.", if Nash-CoT outperforms self-consistency with fewer paths, it indicates that Nash-CoT is better than self-consistency across all settings with an equal or greater number of paths.
Which one performs better between Nash CoT and self-consistency when they have the same number of pathways?
In our ablation experiments to test the impact of the number of pathways on Nash CoT, we found that by adjusting the number of pathways, Nash CoT can surpass self-consistency with fewer pathways on Aqua, CommensenseQA, and Object Tracking. Additionally, in Table 4, we used GLM4 (large), Mistral-Instruct (7B) to outperform self-consistency on the MATH dataset with the same number of pathways.
Appendix H Case Study
H.1 LLM as preference model
In this experiment, we provided examples of LLM inferences to demonstrate that LLMs can select the most appropriate template based on the question. Specifically, in Table 9, we set up a series of templates for different roles, arranging them in different sequences and labeling them with numbers. These were input into the model, allowing it to rank the templates according to its preference for the question. We observe in Table 10 that the order given by the LLM corresponded to the performance strengths of different templates on the respective tasks, which proves that LLMs can directly serve as preference models to select the optimal template for a specific question.
Player Description Correlation
Mathematician You are a mathematician, You excel at analyzing problems from a Mathematical logical perspective and arrive at conclusions that align with your values. High
Student You are a student, please answer the following questions. Medium
Poker Player You are playing a poker game, please do your best to play poker games. Low
Table 9: Player templates we utilized.
Input Ranking
We have four templates: 1. Mathematician: You are a mathematician, You excel at analyzing problems from a Mathematical logical perspective and arrive at conclusions that align with your values. 2. Student: You are a student, please answer the following questions. 3. Poker Player: You are playing a poker game, please do your best to play poker games. Please rank their correlations to a math problem, and directly output the sequential ranked numerical tag. 1, 2, 3
We have four templates: 1. Student: You are a student, please answer the following questions. 2. Mathematician: You are a mathematician, You excel at analyzing problems from a Mathematical logical perspective and arrive at conclusions that align with your values. 3. Poker Player: You are playing poker game, please do your best to play poker games. Please rank their correlations to math problem, and directly output the sequential ranked numerical tag. 2, 1, 3
Table 10: We provide cases for the ranking provided by LLM. The roles of Mathematician, Student, and Poker Player achieved scores of 57.37, 52.45, and 47.54, respectively, on GSM8K. The performance of these different roles is consistent with the model’s ranking.
Meanwhile, we utilize ChatGPT to rank these templates based on their relevance to the topic of questions. Our inputs are shown in Table 10. The model’s ranking of different roles is consistent with their performance levels on the mathematical dataset. Meanwhile, changing the input order of the templates simultaneously does not affect the model’s ranking of the templates.
Appendix I Format of our processed MATH dataset
We provide several cases for readers to understand the format of our processed MATH dataset.
Dataset Question Answer Capacity
Algebra How many vertical asymptotes does the graph of $y=\frac{2}{x^{2}+x-6}$ have? 2 778
Counting and Probability $n$ fair 6-sided dice are simultaneously rolled. The probability that exactly two of them show a number other than 1 is $\frac{25}{216}$. Find $n$. 4 291
Geometry We have triangle $\triangle ABC$ where $AB=AC$ and $AD$ is an altitude. Meanwhile, $E$ is a point on $AC$ such that $AB\parallel DE.$ If $BC=12$ and the area of $\triangle ABC$ is $180,$ what is the area of $ABDE$? 135 235
Intermediate Algebra Find the sum of all complex roots of the equation $\frac{1}{x-1}+\frac{1}{x-5}+\frac{1}{x-10}+\frac{1}{x-25}=2,$ given that there are no repeated roots. 43 443
Number Theory Kirsty needs to hire a plumber to fix her house. The plumber charges $242_{5}$ dollars for every hour of labor and $367_{8}$ dollars for equipment. If the plumber works for $3.5_{10}$ hours, how many dollars (in base ten) will Kirsty owe the plumber? 499 474
Prealgebra John and Gary are playing a game. John spins a spinner numbered with integers from 1 to 20. Gary then writes a list of all of the positive factors of the number spun except for the number itself. Gary then creates a new spinner with all of the numbers on his list. John then spins this spinner, and the process continues. The game is over when the spinner has no numbers on it. If John spins a 20 on his first spin, what is the maximum number of total spins (including the one he already made) that John can make before the game is over? 4 618
Precalculus Let $\mathbf{a}$ and $\mathbf{b}$ be vectors such that $\mathbf{v}=\operatorname{proj}_{\mathbf{a}}\mathbf{v}+\operatorname{proj}_{\mathbf{b}}\mathbf{v}$ for all vectors $\mathbf{v}.$ Enter all possible values of $\mathbf{a}\cdot\mathbf{b},$ separated by commas. 0 133
Table 11: Examples of MATH dataset. We give the format of our processed MATH dataset. Meanwhile, the dataset or its URL will be available on our codebase page upon release.